# Automated Proof Generation for Rust Code via Self-Evolution

### v1

@@ 4.2.4 Impact of Specification Quality
As mentioned in Section 3.2, we only keep specifications with high Correctness scores and reasonably high Completeness scores. To investigate the impact of this design decision, we zoom into the last round of proof-synthesis model fine-tuning to see how the quality of proof-synthesis model can be affected by the quality of specification dataset. Specifically, the current Round-3 SAFE model is obtained by fine-tuning the Round-2 SAFE model using correct proof synthesized by the bootstrapping model, the Round-1 model, and the Round-2 model based on the specifications $S$ selected by SAFE, denoted as $\mathbf{P_{0}[\mathbf{S}]}\cup\mathbf{P_{1}[\mathbf{S}]}\cup\mathbf{P_{2}[\mathbf{S}]}$ ($\mathbf{P}_{i}[\mathbf{S}]$ denotes all the correct proofs synthesized in round $i$ based on specification set $S$). In our two alternative settings, we go back to the set of all specifications synthesized during the spec-synthesis self-evolving procedure and randomly sample a set of specifications $S*$ that have the same size as $S$; naturally $S*$ contains many low-quality specifications that do not pass our Correctness and Completeness threshold. In our alternative setting of ‘Round 2 w/ mix-quality’, we obtain a model by fine-tuning the Round-2 SAFE model using $\mathbf{P_{0}[\mathbf{S}]}\cup\mathbf{P_{1}[\mathbf{S}]}\cup\mathbf{P_{2}[\mathbf{S*}]}$; in our alternative setting of ‘low-quality spec-input’, we obtain a model by fine-tuning the original DeepSeekCoder using $\mathbf{P_{2}[\mathbf{S*}]}$ alone.
From the results in Table 2, we can see that high-quality specifications contribute substantially to the end-to-end effectiveness of SAFE. If the model is trained without any high quality data (‘low-quality spec-input’), the accuracy drops for all three subsets of VerusBench and even drops to 0% Accuracy@1 and 5.26% Accuracy@10 for SV. Probably because more debugging data is created in this setting, the self-debugging feature helps more for this setting, but the overall accuracy still lags way behind the default SAFE Round-3 model. When low quality data is mixed with high quality ones, the performance still drops as shown in the ‘Mixed-Quality’ row in the table. In Appendix, we further report paired t-test results for Table 3 in Table 7, and provide an example Rust program (Listing 14) whose low quality specification led to an extremely simple proof — this simple proof probably offered no benefit for proof-synthesis fine-tuning. In conclusion, specification selection matters to the fine-tuning and the evolution of proof-synthesis models.
@@
Model Metric Proof Generation Self-Debugging
SV MBPP Tutorial Total SV MBPP Tutorial Total
@@
Round 3 Accuracy@10 92.11 35.90 52.17 53.96 97.37 58.97 65.22 70.50
DSCoder Accuracy@1 0.00 29.49 4.35 17.27 0.00 43.59 21.74 28.06
fine-tuned Δ( Round $\Delta(-\mbox{ Round 3})$ (-78.95) (-1.28) (-21.74) (-25.9) (-81.58) (+2.56) (-4.35) (-21.58)
w/ low-quality Accuracy@10 5.26 32.05 21.74 23.02 15.79 50.00 47.83 40.29
spec-input Δ( Round $\Delta(-\mbox{ Round 3})$ (-86.84) (-3.85) (-30.43) (-30.94) (-81.58) (-8.97) (-17.39) (-30.22)
Round 2 Accuracy@1 36.84 30.77 17.39 30.22 42.11 43.59 21.74 39.57
Model Δ( Round $\Delta(-\mbox{ Round 3})$ (-42.11) (-0.00) (-8.7) (-12.95) (-39.47) (+2.56) (-4.35) (-10.07)
fine-tuned Accuracy@10 65.79 32.05 30.43 41.01 71.05 51.28 52.17 56.83
w/ mix-quality Δ( Round $\Delta(-\mbox{ Round 3})$ (-26.32) (-3.85) (-21.74) (-12.95) (-26.32) (-7.69) (-13.04) (-13.67)
@@ 5 Conclusion
In this paper, we propose SAFE, a novel self-evolving framework to advance automated proof generation for Rust. SAFE alleviates the severe data lack challenge by coupling data synthesis and model fine-tuning in a self-evolving manner, demonstrating superior efficiency and precision compared to relying solely on GPT-4o. Through ten of thousands of synthesized proofs and the self-debugging mechanism, we improve the capability of open-source models, initially unacquainted with formal verification, to automatically write proof for Rust code. Our evaluation shows that SAFE leads to a significant improvement over GPT-4o, more than doubling the accuracy of GPT-4o for a human-curated benchmark set and an even bigger advantage on a large LLM-crafted benchmark set.
@@ Appendix A Verus Examples
Imperfect proof example. An imperfect proof annotation of binary search is shown in Listing 2. Compared with the correct version in Listing 1, this imperfect proof lacks a crucial loop invariant in Line 20. According to the error messages following the code, we can see, a postcondition cannot be proved without this loop invariant. Furthermore, the absence of this invariant causes Verus failing to prove the absence of arithmetic underflow/overflows. Specifically, this loop invariant states that value k exists in v[i1 .. i2]. Without this loop invariant, Verus cannot reason about the comparison between v[ix] and k on line 24, and hence cannot reason about how the values of i1 and i2 would change in this loop. This then leads to Verus’ failure in reasoning about the bound of i2 - i1 on line 23 (and hence the arithmetic underflow/overflow error), and the final value of i1 on line 30 (and hence the postcondition not satisfied error).
More complicated proof annotations. Listing 3 shows much more complicated proof annotations for a Rust function that computes the fibonacci sequence. The proof for this code includes not only loop invariants which were introduced in Listing 1, but also a proof function (Line 16–29 in the listing), a proof block (Line 53–55), and assert statements (Line 27) that are used to provide hints to the underlying theorem prover. This example comes from the Verus paper (Lattuada et al., 2023), with more detailed explanation available in that paper.
@@ C.1 Specification Filtering
Listing 7 shows an example of how we leverage test cases and Verus’ symbolic reasoning capability to quickly score and filter specifications. For each function with a synthesized specification (i.e., Line 6–8 in Listing 7, in order to know whether this specification is consistent with a test case $t$, we do the following: 1) we replace the function body with Verus’ assume statements. Each assume tells Verus about the value of one input or output variable (these values come from the test case $t$); 2) when an input/output variable is a container, several assert statements are inserted to make sure the underlying theorem prover can correctly reason about the value of every element in the container; 3) Verus is invoked to prove this function. If the specification is consistent with this test case, verus verification would succeed; otherwise, verus verification would fail.
As mentioned in the paper, to evaluate the completeness of a specification, we mutate existing test cases to see if incorrect test cases can be rejected by a specification. In this example, we mutate the output in the ground truth by adding a new value $15$ into $result$ vector. In this case, Verus will report that the post-condition on Line 7 fails because $15$ does not exist in vector $a$. If a synthesized specification does not contain something like Line 7, this mutated test case will likely point out the incomppleteness of that specification.
In our experiments, we take 5 ground truth test cases and 20 mutated wrong test cases on average for scoring. We filter out all the specifications that have a score lower than 0.8 in correctness or 0.6 in incompleteness. To avoid too many specifications preserved for single program, we allow at most three specifications to be preserved for a single function.
@@ C.2 Statistics of Synthesized Data
As explained earlier in the paper, we obtained 45,495 Rust single-function programs from MBPP training set and CodeNet dataset. After translating them into Verus-compatiable Rust programs, we have 21,398 programs left for specification synthesis. The statistics of all synthesized data in different round of our self-evolution process in shown in Table 4. Note that, at the very beginning, we put aside a random subset of CodeNet Rust programs aside to use for the evaluation of SAFE. Those programs do not participate in the fine-tuning and are not included in the numbers shown in Table 4.
In round 0, we employ GPT-4o to synthesize specification, retaining only the verified Rust functions along with their specifications. This process was time-consuming, taking an entire month to synthesize 3,673 verified programs. Even with our designed prompts, GPT-4o needs to sample 20–30 times for a single program on average. We then use the synthesized specifications to bootstrap the self-evolving specification synthesis process, which is conducted over two rounds. In each round, we leverage the fine-tuned model to generate specifications for all 21k Verus programs and select the high quality ones for further fine-tuning. At the end, we use the selected 19,017 specifications as inputs for the proof synthesis procedure. The numbers of verified programs and debugging fine-tuning triplets are also shown in Table 4. As we can see, as many as 9,706 verified programs and 10,486 debugging data are produced after Round-2 model and are used to fine-tune the Round-3 model, the final proof-synthesis model.
@@ C.3 Training Hyperparameters
Bootstrapped by 3,673 verified programs synthesized by GPT-4o, we employ the DeepSeekCoder model (Guo et al., 2024) as the generator in the self-evolution procedure. We run the self-evolving specification generation and auto-proving process multiple rounds. For specification generation, we run two rounds. Proof generation is three. At each round of fine-tuning, we combine the verified programs and self-debugging data pairs together as training data, and train DeepSeekCoder 5 epochs, using a batch size of 128 and a learning rate of $1\times 10^{-5}$.
Appendix D Additional Experiments
D.1 Statistics of VerusBench
Figure 3 illustrates the token number distributions of specifications and proofs in VerusBench. For preconditions, in Figure 3, we show that SV’s average token number is substantially larger than that of MBPP and Tutorial. In contrast, SV’s average token number is substantially smaller than the rest two as shown in Figure 3. Such statistics can, to some extent, reflect the difficulty of three components in VerusBench, i.e., SV is relatively easiler than the rest two. When comparing the human-written proof annotations for MBPP and Tutorial programs in VerusBench, Figure 3 shows that the average token number of proof annotations in MBPP is larger than that of Tutorial, although there are some outliers.
(a) The token number’s distribution of preconditions
(b) The token number’s distribution of postconditions
(c) The token number’s distribution of ground-truth proofs
@@ D.1 Statistics of VerusBench
D.2 Results of Statistical Significance Testing
@@
Table 5: The p-values of SAFE’s accuracy compared to raw models, non-significant data ($p>0.05$) points are pointed out in Italics
@@
Comparison between Various Rounds of Self-Evolving. Table 6 shows the p-values of SAFE’s accuracy between various rounds of self-evolving. We highlight p-values that are larger than 0.05 in italics. When compared to round 1, round 2 and round 3 show significant improvement in Accuracy, especially after self-debugging. Additionally, we also show that there is insignificant improvement between round 2 and round 3 except the accuracy in Tutorial after self-debugging. Thus, three rounds of self-evolution is sufficient for our approach.
Table 6: The p-values of SAFE’s accuracy between various rounds of self-evolving, non-significant data ($p>0.05$) points are pointed out in Italics
@@
Comparison between SAFE’s Accuracy with Low/High Quality Specifications. Table 7 shows the p-values of SAFE’s accuracy between round 3 and low quality ones. We highlight p-values that are larger than 0.05 in italics. We show that in most scenarios (except Accuracy@1 in round 2 + low quality specifications), low quality specifications might leads to a significant decrease in terms of SAFE’s end-to-end effectiveness.
Table 7: The p-values of SAFE’s accuracy with low quality specifications, compared to round 3 with high quality ones, non-significant data ($p>0.05$) points are pointed out in Italics
@@
DSCoder fine-tuned w/ low-quality spec-input Accuracy@1 5.15E-64 0.79 3.04E-07 3.22E-06 3.62E-71 0.72 0.40 1.98E-04
@@
Round 2 Model fine- tuned w/ mix-quality Accuracy@1 1.94E-13 1.00 8.07E-02 2.45E-02 1.71E-12 0.72 0.40 0.07
@@
D.3 Case Studies
D.3.1 SAFE’s Advantages Compared to GPT-4o

### v2

@@ 4.3.3 Impact of Specification Quality
As mentioned in Section 3.2, we keep only specifications with high Correctness scores and reasonably high Completeness scores. To investigate the impact of this design decision, we zoom into the last round of proof-synthesis model fine-tuning to see how the quality of a proof-synthesis model can be affected by the quality of the used specification dataset. Specifically, the current Round-3 SAFE model is obtained by fine-tuning the Round-2 SAFE model using correct proofs synthesized by the bootstrapping model, the Round-1 model, and the Round-2 model based on the specifications $S$ selected by SAFE, denoted as $\mathbf{P_{0}[\mathbf{S}]}\cup\mathbf{P_{1}[\mathbf{S}]}\cup\mathbf{P_{2}[\mathbf{S}]}$ ($\mathbf{P}_{i}[\mathbf{S}]$ denotes all the correct proofs synthesized in round $i$ based on specification set $S$). In our two alternative settings, we go back to the set of all specifications synthesized during the specification synthesis self-evolving procedure and randomly sample a set of specifications $S*$ that have the same size as $S$; naturally $S*$ contains many low-quality specifications that do not pass our Correctness or Completeness threshold. In our alternative setting of ‘Mix-Quality’, we obtain a model by fine-tuning the Round-2 SAFE model using $\mathbf{P_{0}[\mathbf{S}]}\cup\mathbf{P_{1}[\mathbf{S}]}\cup\mathbf{P_{2}[\mathbf{S*}]}$; in our alternative setting of ‘Low-Quality’, we obtain a model by fine-tuning the original DeepSeekCoder using $\mathbf{P_{2}[\mathbf{S*}]}$ alone.
From the results in Table 3, we can see that high-quality specifications contribute substantially to the end-to-end effectiveness of SAFE. If a model is trained without any high quality data (‘Low-Quality’), the accuracy drops for all three subsets of VerusBench and even drops to 0% Accuracy@1 and 5.26% Accuracy@10 for SV. Probably because more debugging data is created in this setting, the self-debugging feature helps more for this setting, but the overall accuracy still lags way behind the default SAFE Round-3 model. When low quality data is mixed with high quality one, the performance still drops as shown in the ‘Mixed-Quality’ row in the table. In Appendix, we further report paired t-test results for Table 3 in Table 7, and provide an example Rust program (Listing 14) whose low quality specification leads to an extremely simple proof—this simple proof probably offers no benefit for proof-synthesis fine-tuning. In conclusion, specification selection matters to the fine-tuning and the evolution of proof-synthesis models.
@@
Model Metric Proof Generation Self-Debugging ($K+K*K$)
(Spec-Quality) SV MBPP Tutorial Total SV MBPP Tutorial Total
@@
Accuracy@10 92.11 35.90 52.17 53.96 97.37 58.97 65.22 70.50
Low-Quality Accuracy@1 0.00 29.49 4.35 17.27 0.00 43.59 21.74 28.06
Accuracy@10 5.26 32.05 21.74 23.02 15.79 50.00 47.83 40.29
Mix-Quality Accuracy@1 36.84 30.77 17.39 30.22 42.11 43.59 21.74 39.57
Accuracy@10 65.79 32.05 30.43 41.01 71.05 51.28 52.17 56.83
@@ 5 Conclusion
In this paper, we have proposed SAFE, a novel self-evolving framework to advance automated proof generation for Rust. SAFE alleviates the severe data scarcity challenge by coupling data synthesis and model fine-tuning in a self-evolving manner, demonstrating superior efficiency and precision compared to relying solely on GPT-4o. Through ten of thousands of synthesized proofs and the self-debugging mechanism, we improve the capability of open-source models to automatically write proofs for Rust code. Our evaluation shows that SAFE has a significant improvement over GPT-4o.
Acknowledgments
This work was partially supported by National Natural Science Foundation of China under Grant No. 92464301. Tao Xie is also affiliated with the School of Computer Science, Peking University; Key Laboratory of High Confidence Software Technologies (Peking University), Ministry of Education, Beijing, China
@@ Appendix A Verus Examples
A.1 Imperfect Proof Example
An imperfect proof annotation of binary search is shown in Listing 2. Compared with the correct version in Listing 1, this imperfect proof lacks a crucial loop invariant in Line 20. According to the error messages following the code, we can see that a postcondition cannot be proved without this loop invariant. Furthermore, the absence of this invariant causes Verus to fail to prove the absence of arithmetic underflow/overflows. Specifically, this loop invariant states that value k exists in v[i1 .. i2]. Without this loop invariant, Verus cannot reason about the comparison between v[ix] and k on Line 24, and hence cannot reason about how the values of i1 and i2 would change in this loop. This incapability of reasoning then leads to Verus’ failure in reasoning about the bound of i2 - i1 on Line 23 (and hence the arithmetic underflow/overflow error), and the final value of i1 on Line 30 (and hence the postcondition not satisfied error).
A.2 More Complicated Proof Annotations
Listing 3 shows much more complicated proof annotations for a Rust function that computes the Fibonacci sequence. The proof for this code includes not only loop invariants introduced in Listing 1, but also a proof function (Lines 16–29 in the listing), a proof block (Lines 53–55), and assert statements (Line 27) that are used to provide hints to the underlying theorem prover. This example comes from the Verus paper (Lattuada et al., 2023), with more detailed explanation available in that paper.
@@ C.1 Specification Filtering
The example below shows two types of imperfect specification. The specification in the left column is correct but incomplete for the binary search function, as an incorrect implementation that always returns 0 would be accepted by this specification; the specification in the right column below is incorrect, as a correct implementation may fail this specification when the input array has multiple elements matching the search key.
1#A correct, but incomplete specification.
20 <= r,
3r < v.len(),
1#An incorrect specification.
2r > 0 ==> k > v[r - 1],
3r < v.len() - 1 ==> k < v[r + 1],
Listing 7 shows an example of how we leverage test cases and Verus’ symbolic reasoning capability to quickly score and filter specifications. For each function with a synthesized specification (i.e., Lines 6–8 in Listing 7, in order to know whether this specification is consistent with a test case $t$, we do the following: 1) We replace the function body with Verus’ assume statements. Each assume tells Verus about the value of one input or output variable (these values come from the test case $t$); 2) When an input/output variable is a container, multiple assert statements are inserted to make sure that the underlying theorem prover can correctly reason about the value of every element in the container; 3) Verus is invoked to prove this function. If the specification is consistent with this test case, Verus verification would succeed; otherwise, Verus verification would fail.
As mentioned earlier in the paper, to evaluate the completeness of a specification, we mutate existing test cases to see whether incorrect test cases can be rejected by a specification. In this example, we mutate the output in the ground truth by adding a new value $15$ into $result$ vector. In this case, Verus will report that the postcondition on Line 7 fails because $15$ does not exist in vector $a$. If a synthesized specification does not contain something like Line 7, this mutated test case will likely point out the incompleteness of that specification.
In our experiments, we take 5 ground truth test cases and 20 mutated wrong test cases on average for scoring. We filter out all the specifications that have a score lower than 0.8 in correctness or 0.6 in incompleteness. To avoid too many specifications being preserved for a single program, we allow at most three specifications to be preserved for a single function.
@@ C.2 Statistics of Synthesized Data
As explained earlier in the paper, we obtain 45,495 Rust single-function programs from the MBPP training set and CodeNet dataset. After translating them into Verus-compatiable Rust programs, we have 21,398 programs left for specification synthesis. The statistics of all synthesized data in different rounds of our self-evolution process are shown in Table 4. Note that, at the very beginning, we put aside a random subset of CodeNet Rust programs to be used for the evaluation of SAFE. Those programs do not participate in the fine-tuning and are not included in the numbers shown in Table 4.
In round 0, we employ GPT-4o to synthesize specifications, retaining only the verified Rust functions along with their specifications. This process is time-consuming, taking an entire month to synthesize 3,673 verified programs. Even with our designed prompts, GPT-4o needs to sample 20–30 times for a single program on average. We then use the synthesized specifications to bootstrap the self-evolving specification synthesis process, which is conducted over two rounds. In each round, we leverage the fine-tuned model to generate specifications for all 21k Verus programs and select the high quality ones for further fine-tuning. At the end, we use the selected 19,017 specifications as inputs for the proof synthesis procedure. The numbers of verified programs and debugging fine-tuning triplets are also shown in Table 4. As we can see, as many as 9,706 verified programs and 10,486 debugging data are produced after the Round-2 model and are used to fine-tune the Round-3 model, being the final proof-synthesis model.
@@ C.3 Training Hyperparameters
Bootstrapped by 3,673 verified programs synthesized by GPT-4o, we employ the DeepSeekCoder model (Guo et al., 2024) as the generator in the self-evolution procedure. We run the self-evolving specification generation and auto-proving process for multiple rounds. For specification generation, we run two rounds. Proof generation is three. At each round of fine-tuning, we combine the verified programs and self-debugging data pairs together as training data, and train DeepSeekCoder 5 epochs, using a batch size of 128 and a learning rate of $1\times 10^{-5}$.
Appendix D Discussion and Limitation
D.1 Scaling to Real-World Large Software Projects
In this paper, we have focused on generating proofs for small programs mainly because there are too few existing large Verus projects for us to fine-tune LLMs. Consequently, just like many code-synthesis projects that start from small programs, we also focus on small programs as a starting point for Verus proof synthesis.
Since every function is the unit for Verus verification, we believe that the LLM fine-tuned by SAFE on functions in small programs would continue to be useful for functions in large projects. Of course, if we apply SAFE to synthesize proofs for large Rust projects, we expect a key challenge in how to resolve code dependencies across functions. Specifically, a function may call another executable function or specification function, and the callee function may exist in a different file and/or belong to a different class. How to resolve all the code dependencies and provide LLM with all the needed information may require support that goes beyond machine learning.
D.2 Fine-Grained Self-Debugging
Figure 2 illustrates that increasing the number of self-debugging rounds hardly improves proof-synthesis accuracy. This situation might be changed if we change the formation of our self-debugging training data. Currently, the training data for self-debugging includes many pairs of incorrect proof $Y_{X}$ and a corresponding correct proof $Y_{\checkmark}$. Sometimes, the incorrect proof may contain many mistakes, causing many different verification errors. If future research can break down the difference between $Y_{X}$ and $Y_{\checkmark}$, figuring out which edit in $Y_{\checkmark}$ is used to fix which verification error in $Y_{X}$, the resulting data can probably train a model that is better at fixing deeply flawed proof through multiple rounds of debugging.
Appendix E Additional Experiments
E.1 Statistics of VerusBench
Figure 3 illustrates the distributions of the number of tokens in specifications and proofs in VerusBench. For preconditions, in Figure 3, we show that SV’s average number of tokens is substantially larger than that of MBPP and Tutorial. In contrast, SV’s average number of tokens is substantially smaller than the rest two as shown in Figure 3. Such statistics can, to some extent, reflect the difficulty of three components in VerusBench, i.e., SV is relatively easiler than the rest two. When comparing the human-written proof annotations for MBPP and Tutorial programs in VerusBench, Figure 3 shows that the average number of tokens of proof annotations in MBPP is larger than that of Tutorial, although there are some outliers.
(a) The distribution of the number of tokens in preconditions
(b) The distribution of the number of tokens in postconditions
(c) The distribution of the number of tokens in ground-truth proofs
@@ E.1 Statistics of VerusBench
E.2 Distribution of Specification Metrics
Figure 4 shows the distribution of the correctness-score and the completeness-score of all the specifications synthesized during the self-evolving process of SAFE. As we can see, during the self-evolving process, many trivially correct specifications are generated (i.e., $1.0$ correctness score but $<0.6$ completeness score)—the clusters at the top of the figure and not overlapping with the gray rectangle. These specifications tend to lead to trivial proofs and hence are filtered out. There is also a big cluster of incorrect specifications that have incorrectly rejected all test cases at the bottom-right corner of the figure (i.e., 00 correctness score but $1.0$ completeness score). There is no way Verus can prove that the Rust function satisfies these specifications, no matter what proof annotations are used, so it is good that SAFE filters all of them out.
Figure 4: The correctness and completeness distribution of SAFE synthesized specifications. The coordinate of the center of every circle indicates the Completeness-score (x-axis) and the Correctness-score (y-axis) of the corresponding cluster of specifications; the size and color of each circle reflect the number of specifications that are in one cluster. The gray rectangle highlights the range of accepted scores: all the specification clusters that overlap with the gray rectangle are considered good enough and hence kept by SAFE.
E.3 Results of Statistical Significance Testing
@@
Table 5: The p-values of SAFE’s accuracy compared to raw models; non-significant data ($p>0.05$) points are pointed out in Italics
@@
Comparison between Various Rounds of Self-Evolving. Table 6 shows the p-values of SAFE’s accuracy between various rounds of self-evolving. We highlight p-values that are larger than 0.05 in italics. When compared to round 1, round 2 and round 3 show significant improvement in Accuracy, especially after self-debugging. Additionally, we also show that there is insignificant improvement between round 2 and round 3 except the accuracy in Tutorial after self-debugging. Thus, three rounds of self-evolution are sufficient for our approach.
Table 6: The p-values of SAFE’s accuracy between various rounds of self-evolving; non-significant data ($p>0.05$) points are pointed out in Italics
@@
Comparison between SAFE’s Accuracy with Low/High Quality Specifications. Table 7 shows the p-values of SAFE’s accuracy between round 3 and low quality ones. We highlight p-values that are larger than 0.05 in italics. We show that in most scenarios (except Accuracy@1 in round 2 + low quality specifications), low quality specifications might lead to a significant decrease in terms of SAFE’s end-to-end effectiveness.
Table 7: The p-values of SAFE’s accuracy with low quality specifications, compared to round 3 with high quality ones; non-significant data ($p>0.05$) points are pointed out in Italics
@@
Low-Quality Accuracy@1 5.15E-64 0.79 3.04E-07 3.22E-06 3.62E-71 0.72 0.40 1.98E-04
@@
Mix-Quality Accuracy@1 1.94E-13 1.00 8.07E-02 2.45E-02 1.71E-12 0.72 0.40 0.07
@@
E.4 Self-Evolution with Small Backbone Models
To evaluate the effectiveness of our self-evolution framework on LLMs with a smaller number of parameters, we conduct additional experiments using DeepSeekCoder-1.3B (DSCoder-1.3B) as the backbone model. In this experiment, we follow the same experimental settings described in Section 4.1, bootstrapping DSCoder-1.3B with GPT-4o’s predictions.
After the same rounds of self-evolving specification generation and proof generation, the results on VerusBench are listed in Table 8. The results indicate that our self-evolution framework effectively improves the accuracy even with a smaller backbone model of 1.3 billion parameters.
Table 8: Each round’s accuracy with DeepSeekCoder-1.3B as the backbone model
Proof Generation Self Debugging ($k+k*k$)
Metric Accuracy@1 Accuracy@10 Accuracy@1 Accuracy@10
Raw 1.44 6.47 - -
Round 1 12.95 24.46 12.95 24.46
Round 2 19.42 26.69 24.46 52.52
Round 3 21.58 40.29 27.34 57.55
Appendix F Case Studies
F.1 SAFE’s Advantages Compared to GPT-4o